OpenStudy (anonymous):

Sally has seven times as many dimes as nickels. Their value is $3.00. How many nickels does she have?

OpenStudy (anonymous):

0.75n= 3

OpenStudy (anonymous):

No giving Answers

OpenStudy (anonymous):

but u did

OpenStudy (anonymous):

no i didn't

OpenStudy (anonymous):

stop staring fights in a question

OpenStudy (anonymous):

First of all @kathy0514 : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Let d and n be the number of dimes and nickles respectively. Express the problem using pennies. From the problem statement the following is true. \[\left\{\frac{d}{n}=7,10d+5n=300\right\} \]

OpenStudy (anonymous):

FYI I am not new and last time stop starting fights in chat or I will call ash